OpenStudy (anonymous):

RIGHT TRIANGLE... opposite=2x, adjacent=x+1, and hypotenuse=x+19...find x THA ANSWERS SUPPOSED TO BE 15 BUT IDK HOW TO GET THERE. :'(

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

hmm... so how do you know is .. .15 though?

OpenStudy (jdoe0001):

ohh.. you saw it on the back of the book... I see

OpenStudy (anonymous):

WELL YEAH, KINDA...THE ETACHER GAVE US THE ANSWERS...BUT YOU HAVE TO SHOW WORK, *oops it was on upper case

OpenStudy (anonymous):

youre supposed to factor using the pythagorean theorem

OpenStudy (jdoe0001):

soo... ok... how far down have you tried?

OpenStudy (anonymous):

so far i have -13x=square root of 180

OpenStudy (anonymous):

but idk if thats right

OpenStudy (anonymous):

By the pythagorean theorem: (2x)^2 + (x+1)^2 = (x+19)^2 => 4x^2 + x^2 + 2x + 1 = x^2 + 38x + 361 => 4x^2 - 36x - 360 = 0 --> factor out 4 => x^2 - 9x - 90 = 0 => (x-15)(x+6) = 0 => x = 15 or x = -6 Since 'x' has to be a positive number because it's involved in the lengths of the sides of the triangles which means that we can discard the solution x = -6. Therefore x = 15. @merylzac